# Multilayer network science: theory, methods, and applications

### v1

@@ II.3.2 Higher-order interactions
Group interactions are widespread in real-world systems, ranging from social dynamics Iacopini2019; Arruda2020; Battiston2021; Landry2020; St-Onge2022; Arruda2023; Kiss2023, such as group chats, to chemical reactions involving multiple reactants Jost2019; Battiston2021; Traversa2023. In these contexts, the traditional network science framework, which considers only pairwise interactions, is often insufficient to capture the complex dependencies between multiple interacting entities. Higher-order models provide a richer framework for accurately representing these systems and avoiding misleading predictions Battiston2020; Battiston2021; Arruda2023; FerrazdeArruda2024Aug. For example, contagion processes that model the population as a graph will present a second-order phase transition Pastor-Satorras2015; Arruda2018. However, contagion models in higher-order networks often exhibit discontinuous transitions Iacopini2019; Arruda2020; Landry2020; St-Onge2022; Arruda2023; FerrazdeArruda2024Aug. Similarly, in pairwise contagion models, one expects a single stable solution, while multistability is possible in contagion models in higher-order networks Arruda2023; Kiss2023; FerrazdeArruda2024Aug. Interestingly, these new phenomena are not restricted to contagion processes; they have also been observed in synchronization processes Skardal2023. These examples demonstrate how neglecting the higher-order nature of a process can alter the outcome of a model, potentially leading to incorrect predictions.
@@
Mathematically, higher-order interactions are usually modelled using hypergraphs or simplicial complexes Battiston2020; Boccaletti2023; Torres2021. Hypergraphs allow edges (hyperedges) to connect any number of nodes, whereas simplicial complexes require that all subsets of a hyperedge be present. The best choice of representation often depends on the dynamics being studied. A discussion about this topic can be found in Torres2021.
@@
Figure 5: Schematic representation of different networked systems. The bottom row depicts systems with pairwise interactions, while the top row contains group interactions including more than two nodes. As with classical pairwise models, it is possible to associate a type to each interaction, extending the framework of multilayer networks to multilayer hypergraphs. Image adapted from FerrazdeArruda2024Aug.
@@
Compared to pairwise networks, the structural and spectral analysis of higher-order systems is still in development. While concepts such as adjacency and Laplacians are well-defined in networks, their hypergraph counterparts are more complex and often combine structural and dynamical parameters Schaub2020; Mulas2020; Arruda2021. This complicates the analytical study of processes such as diffusion and synchronization, and different definitions may yield different outcomes depending on the context. Similarly, most inference methods in the literature deal with inferring pairwise relationships. Detecting higher-order structures requires new methods Peixoto2020; Contisciani2022.
@@
Recent work has begun to explore multilayer extensions of higher-order systems (Fig. 5), such as triadic percolation processes and interdependent dynamics across layers Sun2023; Chang2023. In the context of spreading processes, a two-layer approach considering the concurrent spread of information and disease was studied in fan_epidemics_2022; hong_coupled_2023. Furthermore, the diffusion of resources coupled with epidemic spreading was examined in  sun_diffusion_2022. These studies demonstrate how combining multilayer and higher-order structures can lead to novel, emergent behaviours, such as oscillatory or chaotic dynamics. Another promising research direction is studying temporal hypergraphs, which extend these representations to time-varying settings. This framework captures interactions involving groups of nodes that occur over finite durations. Examples include simultaneous face-to-face contacts di_gaetano_percolation_2024 and group-level dynamics in social settings Iacopini2024Aug. Temporal hypergraphs provide a natural language to describe and model these evolving group structures, opening new avenues for analysing their evolution and how they influence dynamical processes.
@@ II.3.3 Machine learning
Graph representation learning, or graph embedding, aims to map nodes into a geometric space such that structural properties are preserved goyal18; cai18. This task is essential for applying machine learning techniques to relational data and has been widely used for link prediction, node classification, routing, and visualization. In the past few years, traditional embedding methods have been extended to deal with heterogeneous graphs and knowledge graphs, where nodes or edges carry additional categorical, numerical or textual information. Multiplex networks can be seen as a specific type of heterogeneous graph and have prompted the development of dedicated embedding techniques.
@@
A good multiplex embedding method should effectively capture both intralayer and interlayer structures. Multiplex network embeddings can be divided into two major classes: joint embeddings, which project all replicas of a node across layers to a single point in a low-dimensional Euclidean space, and individual embeddings, where each layer has its own representation of the node. Joint embeddings can be obtained by merging individual embeddings. The general approach followed by most multiplex embedding techniques consists of three phases: sampling, training, and optimization.
@@
Figure 6: General approach to multiplex embedding. Given a multiplex network, the first step involves identifying node pairs that are supposed to have similar (positive samples) and dissimilar (negative samples) representations. Next, the learning framework is used to train the node representations based on the sampled data. The node representations can be learnt for the entire multiplex (bottom right) or for individual layers (top right). Individual embeddings may be fused together to obtain joint embeddings. Courtesy of Siddharth Patwardhan.
@@
Sampling. While embedding, we want pairs of nodes that are close to each other in the network to also have similar embeddings. These are called “positive node pairs” and can be found using several strategies. Techniques such as DeepWalk perozzi14 and node2vec grover16 simulate random walks to define contexts, akin to word embeddings in natural language processing mikolov13. These approaches have been extended to multiplex networks by allowing walks to transition between layers with some probability liu17. Other methods define proximity using K-nearest neighbors or neighborhood-based criteria, as in GraphSAGE hamilton17. To reduce computational cost, negative sampling is used to approximate objective functions, drawing random “negative pairs” assumed to be dissimilar cai18.
@@
Training. In this phase, multiplex embedding methods learn representations using shallow neural networks, although more advanced architectures have been proposed. In shallow neural networks, the training process is specified by the structural organization of the learning algorithm, which can be layerwise or collective learning (Fig. 6, top and bottom rows respectively). Layerwise learning separately trains node embeddings for each layer and combines them using weights, often learned via attention mechanisms qu17; xie20. Collective learning directly learns unified representations by leveraging multiplex sampling and skipgram-style optimizations zhang18; liu17. While shallow networks are common, they face important limitations, such as the inability to generalize to unseen nodes or share parameters between nodes (since the encoder optimizes a unique embedding vector for each node). To alleviate these limitations, more sophisticated learning architectures, like graph neural networks, have been proposed.
@@
Optimization. Embeddings are adjusted to minimize loss functions that reward proximity between positive pairs and penalize proximity between negative ones. Some methods also include node attributes to improve representation quality feng19; cen19. These optimized embeddings can then be used in downstream tasks such as:
@@
  • Link prediction: predicting missing or future links within a layer (layer-wise), across layers (interlayer) or in an aggregated view of the multilayer liu17; xie20; zhang18; feng19; gallo24.
@@
  • Node classification: predicting labels for unlabelled nodes based on known node attributes or partial label data zhang18.
@@
  • Community detection: discovering modular structures across layers by learning weights that reflect each layer’s contribution jing21; cai21.
@@
In addition to Euclidean embeddings, hyperbolic embeddings provide a geometry better suited to hierarchical and navigable networks garcia19. These approaches can potentially improve performance in tasks like greedy routing, where shortest paths are found using only local information. Recent work using layerwise hyperbolic embeddings suggests that interlayer geometric correlations improve routing performance kleineberg16. Future work could explore hyperbolic multiplex embeddings more systematically.
@@
Lastly, synthetic multiplex models bazzi20 can be used to systematically vary layer-specific properties and interlayer correlations, providing testbeds for evaluating embedding performance under controlled conditions. These models support the study of how properties like modularity and community alignment across layers affect embedding quality and task performance.
@@ III.1 Interdependent systems
The importance of multilayer interdependent networks was first recognized by engineers and later formalized through the interdependent percolation model introduced in buldyrev_2010. In addition to standard intralayer links, these systems feature dependency links that connect nodes across layers, such that the functionality of a node in one network depends on the functionality of a node in another. This structure can lead to cascading failures as a breakdown in one layer propagates to others, potentially causing large-scale system collapse. Unlike standard percolation, where failure progresses gradually, interdependent networks can experience abrupt, catastrophic transitions. Notably, networks with heterogeneous degree distributions, typically robust against random failures, become vulnerable in the presence of dependency links, highlighting important design and resilience considerations.
Insights into these phenomena have been obtained through both analytical and numerical studies gao2012networks; Havlin2015Mar; Radicchi2015Jul. In interdependent random networks, the percolation transition is of a hybrid nature: it is discontinuous but also exhibits critical phenomena, such as the presence of critical exponents Baxter2012Dec. In plain percolation, there are two equivalent theoretical descriptions, one based on cluster statistics and the other on the order parameter stauffer92. Yet, these approaches split in the interdependent case, leading to two sets of critical exponents related by scaling relations lee16; choi24. In a recent paper, an interesting relationship between the hybrid transition in cascading processes and the spinodal one was shown, leading to the identification of universality classes Bonamassa2025Feb.
One of the most exciting recent developments is the experimental realization of interdependent percolation in coupled superconducting multilayer systems Bonamassa2023Aug. In this setup, two layers of disordered superconductors are separated by an insulating layer but become effectively coupled through Joule heat. This way, when one superconducting grain transitions to a normal state, it generates heat that causes other grains in the adjacent layer to transition as well. The resulting feedback loop drives a non-local dependency coupling and confirms the theoretical prediction of an abrupt transition. These experimental observations not only validate the theory but also open the door to innovative technological applications based on interdependent dynamics.
@@ III.2 Spreading processes
In the context of diffusion and spreading dynamics, multilayer networks offer a powerful framework to study how coupled processes may interact, leading to the uncovering of hidden dependencies, feedback loops, and emergent phenomena deArruda2018Oct. This approach is particularly valuable when different types of interactions or dynamical processes influence each other in ways that are not evident in single-layer representations.
The range of systems that benefit from this approach is broad. For example, the mutual influence between public opinion and disease spread can be modelled by layering the dynamics of information diffusion with those of epidemic propagation. Such an approach reveals insights into behavioural responses and feedback mechanisms that would otherwise be obscured Granell2013Sep; Wang2015Dec; Fosch2023Nov; peng_multilayer_2021. Multilayer networks also naturally represent systems involving multiple domains with interdependent effects, such as the bidirectional impact between pandemics and economic activity (Fig. 7Pangallo2024Feb. Or systems with a single process but with different spreading properties depending on the type of interaction Aleta2020Jul.
Another compelling example involves the interplay between diseases with distinct transmission mechanisms. Tuberculosis (TB), transmitted via airborne particles, and HIV, spread through sexual contact, represent independent but coexisting threats within a population. Using a multilayer framework, one can explore how interventions or behavioural changes in one disease layer affect the dynamics of the other Sanz2014Oct; Bosetti2020Dec; Kinsley2020Sep; Wang2019Sep. This holistic modelling approach aids in designing integrated public health strategies.
@@
Figure 7: Schematic representation of an epidemic-economic model. The figure illustrates the interplay between two very different dynamics. An economic component, which represents the flow of goods and services among industries and to final consumers through an input-output network, and an epidemic component, tracking pathogen exposure in various settings such as workplaces, community/consumption venues, schools, and households via a multilayer contact network. Agents within the system exhibit diverse socio-economic characteristics (see the accompanying box). The tight integration of the economic and epidemic modules is evident: the epidemic module influences economic outcomes by reducing consumption due to infection concerns, while the economic module affects epidemic transmission by modifying workplace and community interactions through industry-specific employment fluctuations Pangallo2024Feb
@@ III.3 Computational social science
Social systems are inherently characterized by multiple types of relationships that coexist and interact across diverse contexts, timescales, and communication channels. In computational social science, multilayer networks have become a key framework for analysing how behaviours, attitudes, and information propagate through these interconnected pathways. This approach enables the study of complex dynamics such as the interplay between online and offline ties hristova2014keep, cross-platform relationship evolution urena2024multiplexity, and behaviours emerging from interacting relationship types garcia2015ideological.
@@
In political communication, distinct interaction types (support, comments, likes) form separate layers of a multiplex network, each reflecting unique dimensions of polarization and temporal evolution garcia2015ideological. For example, in parliamentary Twitter (now X) networks, echo chambers differ across layers: following networks are highly polarized, while mention networks allow more cross-ideological communication del2018echo. Cross-platform studies find consistent drivers of polarization, such as homophily and selective exposure, even when patterns vary between platforms cinelli2021echo. In a study of French Twitter users, the multidimensional character of polarization was revealed by unfolding the interaction between the layer of professional politicians and that of ordinary users peralta24.
@@
Temporal multilayer networks reveal that social multiplexity itself is dynamic. Relationship types activate in different contexts, and temporal correlations between layers modify how influence and information spread urena2024multiplexity. Human activity’s bursty nature can create both real and spurious correlations starnini2017effects. In signed networks, research has explored how structural balance and polarization evolve, including dynamical equilibrium where macro-level patterns persist despite micro-level changes talaga2023polarization; gonzalez2025evidence.
@@
In the semantic domain, multilayer networks help uncover how communities structure discourse around complex issues. Modelling words, themes, and temporal evolution as distinct layers reveals patterns of resilience and vulnerability in discussions, such as those following disasters randazzo2024multilayer. Temporal text networks vega2018foundations and multimodal network models bonifazi2023multilayer further support analysis of communication across media and platforms.
@@
Multilayer approaches have also enriched opinion dynamics, where single-layer models often oversimplify how ideas form and spread. This approach provides a powerful framework for exploring how phenomena in one domain can influence behaviour in another. For example, a model using a communication layer for sharing opinions and an influence layer for observing actions, revealed a range of complex real-world outcomes zino_two-layer_2020. These included rapid, widespread paradigm shifts, the establishment of unpopular norms where innovation adoption fails despite positive sentiment, and persistent community resistance to change.
@@
As discussed in the previous section, the interaction between different processes is a key area of study for multilayer network research. A recent work amato_opinion_2017 investigated a two-layered network, where agents held opinions on different issues in each layer. While each layer individually tended toward consensus, the interactions between layers could stabilize a mixed state where both opinions coexisted for a long period. This was most pronounced in sparse, positively correlated networks, which facilitated the formation of cross-layer opinion groups.
@@
The adaptive voter model holme06 has been extended to multilayer networks to study how evolving connections affect collective opinion. While multiplexity (the fraction of common nodes across layers) can prevent complete fragmentation, it can also give rise to a new “shattered” fragmentation phase diakonova_irreducibility_2016. This phase is characterized by large consensus groups of opposite opinions coexisting with isolated nodes. A similar shattered fragmentation transition in the multiplex adaptive voter model with a notion of triadic closure has been used to describe the empirical distribution of cluster sizes in an online game klimek2016dynamical.
@@
Furthermore, a central question in this modelling paradigm is the extent to which such complex multilayer systems can be simplified, especially whether the dynamics on a multilayer network could be reduced to an equivalent single-layer representation. It turns out that conventional aggregation procedures fail to capture key non-linear effects, such as the prolonged lifetime of multiplex states at low interlayer connectivity diakonova_irreducibility_2016. This work suggests that there are fundamental, non-reducible differences between multilayer dynamics and their single-layer approximations, highlighting the necessity of explicitly modelling these intricate, interconnected structures to fully understand complex social phenomena.

### v2

@@ II.3.2 Higher-order interactions
Group interactions are widespread in real-world systems, ranging from social dynamics Iacopini et al. (2019); de Arruda, Guilherme Ferraz et al. (2020); Battiston et al. (2021); Landry and Restrepo (2020); St-Onge et al. (2022); Ferraz de Arruda et al. (2023); Kiss et al. (2023), such as group chats, to chemical reactions involving multiple reactants Jost and Mulas (2019); Battiston et al. (2021); Traversa et al. (2023). In these contexts, the traditional network science framework, which considers only pairwise interactions, is often insufficient to capture the complex dependencies between multiple interacting entities. Higher-order models provide a richer framework for accurately representing these systems and avoiding misleading predictions Battiston et al. (2020, 2021); Ferraz de Arruda et al. (2023, 2024). For example, contagion processes that model the population as a graph will present a second-order phase transition Pastor-Satorras et al. (2015); de Arruda et al. (2018). However, contagion models in higher-order networks often exhibit discontinuous transitions Iacopini et al. (2019); de Arruda, Guilherme Ferraz et al. (2020); Landry and Restrepo (2020); St-Onge et al. (2022); Ferraz de Arruda et al. (2023, 2024). Similarly, in pairwise contagion models, one expects a single stable solution, while multistability is possible in contagion models in higher-order networks Ferraz de Arruda et al. (2023); Kiss et al. (2023); Ferraz de Arruda et al. (2024). Interestingly, these new phenomena are not restricted to contagion processes; they have also been observed in synchronisation processes Skardal et al. (2023) and evolutionary games Civilini et al. (2024). These examples demonstrate how neglecting the higher-order nature of a process can alter the outcome of a model, potentially leading to incorrect predictions.
@@
Mathematically, higher-order interactions are usually modelled using hypergraphs or simplicial complexes Battiston et al. (2020); Boccaletti et al. (2023); Torres et al. (2021). Hypergraphs allow edges (hyperedges) to connect any number of nodes, whereas simplicial complexes require that all subsets of a hyperedge be present. The best choice of representation often depends on the dynamics being studied. A discussion about this topic can be found in Torres et al. (2021).
@@
Figure 5: Schematic representation of different networked systems. The bottom row depicts systems with pairwise interactions, while the top row contains group interactions including more than two nodes. As with classical pairwise models, it is possible to associate a type to each interaction, extending the framework of multilayer networks to multilayer hypergraphs. Adapted figure with permission from Ferraz de Arruda et al. (2024). ©2024 by Springer Nature.
@@
Compared to pairwise networks, the structural and spectral analysis of higher-order systems is still in development. While concepts such as adjacency and Laplacians are well-defined in networks, their hypergraph counterparts are more complex and often combine structural and dynamical parameters Schaub et al. (2020); Mulas et al. (2020); Ferraz de Arruda et al. (2021). This complicates the analytical study of processes such as diffusion and synchronisation, and different definitions may yield different outcomes depending on the context. Similarly, most inference methods in the literature deal with inferring pairwise relationships. Detecting higher-order structures requires new methods Young et al. (2021); Contisciani et al. (2022).
@@
Recent work has begun to explore multilayer extensions of higher-order systems (Fig. 5), such as triadic percolation processes and interdependent dynamics across layers Sun et al. (2023); Chang et al. (2023). In the context of spreading processes, a two-layer approach considering the concurrent spread of information and disease was studied in Fan et al. (2022); Hong et al. (2023). Furthermore, the diffusion of resources coupled with epidemic spreading was examined in  Sun et al. (2022). These studies demonstrate how combining multilayer and higher-order structures can lead to novel, emergent behaviours, such as oscillatory or chaotic dynamics. Another promising research direction is studying temporal hypergraphs, which extend these representations to time-varying settings. This framework captures interactions involving groups of nodes that occur over finite durations. Examples include simultaneous face-to-face contacts Di Gaetano et al. (2024); Gallo et al. (2024a) and group-level dynamics in social settings Iacopini et al. (2024). Temporal hypergraphs provide a natural language to describe and model these evolving group structures, opening new avenues for analysing their evolution and how they influence dynamical processes.
@@ II.3.3 Machine learning
Graph representation learning, or graph embedding, aims to map nodes into a geometric space such that structural properties are preserved Goyal and Ferrara (2018); Cai et al. (2018). This task is essential for applying machine learning techniques to relational data and has been widely used for link prediction, node classification, routing, and visualisation. In the past few years, traditional embedding methods have been extended to deal with heterogeneous graphs and knowledge graphs, where nodes or edges carry additional categorical, numerical or textual information Shi et al. (2022). Multiplex networks can be seen as a specific type of heterogeneous graph and have prompted the development of dedicated embedding techniques.
@@
A good multiplex embedding method should effectively capture both intralayer and interlayer structures. Multiplex network embeddings can be divided into two major classes: joint embeddings, which project all replicas of a node across layers to a single point in a low-dimensional Euclidean space, and individual embeddings, where each layer has its own representation of the node. Joint embeddings can be obtained by merging individual embeddings. The general approach followed by most multiplex embedding techniques consists of three phases: sampling, training, and optimisation.
@@
Figure 6: General approach to multiplex embedding. Given a multiplex network, the first step involves identifying node pairs that are supposed to have similar (positive samples) and dissimilar (negative samples) representations. Next, the learning framework is used to train the node representations based on the sampled data. The node representations can be learnt for the entire multiplex (bottom right) or for individual layers (top right). Individual embeddings may be fused together to obtain joint embeddings.
@@
Sampling. While embedding, we want pairs of nodes that are close to each other in the network to also have similar embeddings. These are called “positive node pairs” and can be found using several strategies. Techniques such as DeepWalk Perozzi et al. (2014) and node2vec Grover and Leskovec (2016) simulate random walks to define contexts, akin to word embeddings in natural language processing Mikolov (2013). These approaches have been extended to multiplex networks by allowing walks to transition between layers with some probability Liu et al. (2017). Other methods define proximity using K-nearest neighbors or neighborhood-based criteria, as in GraphSAGE Hamilton et al. (2017). To reduce computational cost, negative sampling is used to approximate objective functions, drawing random “negative pairs” assumed to be dissimilar Cai et al. (2018).
@@
Training. In this phase, multiplex embedding methods learn representations using shallow neural networks, although more advanced architectures have been proposed. In shallow neural networks, the training process is specified by the structural organisation of the learning algorithm, which can be layerwise or collective learning (Fig. 6, top and bottom rows respectively). Layerwise learning separately trains node embeddings for each layer and combines them using weights, often learned via attention mechanisms Qu et al. (2017); Xie et al. (2020). Collective learning directly learns unified representations by leveraging multiplex sampling and skipgram-style optimisations Zhang et al. (2018); Liu et al. (2017). While shallow networks are common, they face important limitations, such as the inability to generalize to unseen nodes or share parameters between nodes (since the encoder optimises a unique embedding vector for each node). To alleviate these limitations, more sophisticated learning architectures, like graph neural networks, have been proposed.
@@
Optimisation. Embeddings are adjusted to minimize loss functions that reward proximity between positive pairs and penalize proximity between negative ones. Some methods also include node attributes to improve representation quality Feng et al. (2019); Cen et al. (2019). These optimised embeddings can then be used in downstream tasks such as:
@@
  • Link prediction: predicting missing or future links within a layer (layer-wise), across layers (interlayer) or in an aggregated view of the multilayer Liu et al. (2017); Xie et al. (2020); Zhang et al. (2018); Feng et al. (2019); Gallo et al. (2024b).
@@
  • Node classification: predicting labels for unlabelled nodes based on known node attributes or partial label data Zhang et al. (2018).
@@
  • Community detection: discovering modular structures across layers by learning weights that reflect each layer’s contribution Jing et al. (2021); Cai et al. (2021).
@@
In addition to Euclidean embeddings, hyperbolic embeddings provide a geometry better suited to hierarchical and navigable networks García-Pérez et al. (2019). These approaches can potentially improve performance in tasks like greedy routing, where shortest paths are found using only local information. Recent work using layerwise hyperbolic embeddings suggests that interlayer geometric correlations improve routing performance Kleineberg et al. (2016). Future work could explore hyperbolic multiplex embeddings more systematically.
@@
Lastly, synthetic multiplex models Bazzi et al. (2020) can be used to systematically vary layer-specific properties and interlayer correlations, providing testbeds for evaluating embedding performance under controlled conditions. These models support the study of how properties like modularity and community alignment across layers affect embedding quality and task performance.
@@ III.1 Interdependent systems
The importance of multilayer interdependent networks was first recognized by engineers and later formalized through the interdependent percolation model introduced in Buldyrev et al. (2010). In addition to standard intralayer links, these systems feature dependency links that connect nodes across layers, such that the functionality of a node in one network depends on the functionality of a node in another. This structure can lead to cascading failures as a breakdown in one layer propagates to others, potentially causing large-scale system collapse. Unlike standard percolation, where failure progresses gradually, interdependent networks can experience abrupt, catastrophic transitions. Notably, networks with heterogeneous degree distributions, typically robust against random failures, become vulnerable in the presence of dependency links, highlighting important design and resilience considerations.
Insights into these phenomena have been obtained through both analytical and numerical studies Gao et al. (2012); Havlin et al. (2015); Radicchi (2015). In interdependent random networks, the percolation transition is of a hybrid nature: it is discontinuous but also exhibits critical phenomena, such as the presence of critical exponents Baxter et al. (2012). In plain percolation, there are two equivalent theoretical descriptions, one based on cluster statistics and the other on the order parameter Stauffer and Aharony (2018). Yet, these approaches split in the interdependent case, leading to two sets of critical exponents related by scaling relations Lee et al. (2016); Choi et al. (2024). In a recent paper, an interesting relationship between the hybrid transition in cascading processes and the spinodal one was shown, leading to the identification of universality classes Bonamassa et al. (2025).
One of the most exciting recent developments is the experimental realization of interdependent percolation in coupled superconducting multilayer systems Bonamassa et al. (2023). In this setup, two layers of disordered superconductors are separated by an insulating layer but become effectively coupled through Joule heat. This way, when one superconducting grain transitions to a normal state, it generates heat that causes other grains in the adjacent layer to transition as well. The resulting feedback loop drives a non-local dependency coupling and confirms the theoretical prediction of an abrupt transition. These experimental observations not only validate the theory but also open the door to innovative technological applications based on interdependent dynamics.
@@ III.2 Spreading processes
In the context of diffusion and spreading dynamics, multilayer networks offer a powerful framework to study how coupled processes may interact, leading to the uncovering of hidden dependencies, feedback loops, and emergent phenomena de Arruda et al. (2018). This approach is particularly valuable when different types of interactions or dynamical processes influence each other in ways that are not evident in single-layer representations.
The range of systems that benefit from this approach is broad. For example, the mutual influence between public opinion and disease spread can be modelled by layering the dynamics of information diffusion with those of epidemic propagation. Such an approach reveals insights into behavioural responses and feedback mechanisms that would otherwise be obscured Granell et al. (2013); Wang et al. (2015); Fosch et al. (2023); Peng et al. (2021). Multilayer networks also naturally represent systems involving multiple domains with interdependent effects, such as the bidirectional impact between pandemics and economic activity (Fig. 7Pangallo et al. (2024). Or systems with a single process but with different spreading properties depending on the type of interaction Aleta et al. (2020a).
Another compelling example involves the interplay between diseases with distinct transmission mechanisms. Tuberculosis (TB), transmitted via airborne particles, and HIV, spread through sexual contact, represent independent but coexisting threats within a population. Using a multilayer framework, one can explore how interventions or behavioural changes in one disease layer affect the dynamics of the other Sanz et al. (2014); Bosetti et al. (2020); Kinsley et al. (2020); Wang et al. (2019). This holistic modelling approach aids in designing integrated public health strategies.
@@
Figure 7: Schematic representation of an epidemic-economic model. The figure illustrates the interplay between two very different dynamics. An economic component, which represents the flow of goods and services among industries and to final consumers through an input-output network, and an epidemic component, tracking pathogen exposure in various settings such as workplaces, community/consumption venues, schools, and households via a multilayer contact network. Agents within the system exhibit diverse socio-economic characteristics (see the accompanying box). The tight integration of the economic and epidemic modules is evident: the epidemic module influences economic outcomes by reducing consumption due to infection concerns, while the economic module affects epidemic transmission by modifying workplace and community interactions through industry-specific employment fluctuations. Reprinted figure with permission from Pangallo et al. (2024). ©2024 by Springer Nature.
@@ III.3 Computational social science
Social systems are inherently characterised by multiple types of relationships that coexist and interact across diverse contexts, timescales, and communication channels. In computational social science, multilayer networks have become a key framework for analysing how behaviours, attitudes, and information propagate through these interconnected pathways. This approach enables the study of complex dynamics such as the interplay between online and offline ties Hristova et al. (2014), cross-platform relationship evolution Ureña-Carrión et al. (2024), and behaviours emerging from interacting relationship types Garcia et al. (2015).
@@
In political communication, distinct interaction types (support, comments, likes) form separate layers of a multiplex network, each reflecting unique dimensions of polarization and temporal evolution Garcia et al. (2015). For example, in parliamentary Twitter (now X) networks, echo chambers differ across layers: following networks are highly polarized, while mention networks allow more cross-ideological communication Del Valle and Bravo (2018). Cross-platform studies find consistent drivers of polarization, such as homophily and selective exposure, even when patterns vary between platforms Cinelli et al. (2021). In a study of French Twitter users, the multidimensional character of polarization was revealed by unfolding the interaction between the layer of professional politicians and that of ordinary users Peralta et al. (2024).
@@
Temporal multilayer networks reveal that social multiplexity itself is dynamic. Relationship types activate in different contexts, and temporal correlations between layers modify how influence and information spread Ureña-Carrión et al. (2024). Human activity’s bursty nature can create both real and spurious correlations Starnini et al. (2017). In signed networks Diaz-Diaz et al. (2025), research has explored how structural balance and polarization evolve, including dynamical equilibrium where macro-level patterns persist despite micro-level changes Talaga et al. (2023); González-Casado et al. (2025).
@@
In the semantic domain, multilayer networks help uncover how communities structure discourse around complex issues. Modelling words, themes, and temporal evolution as distinct layers reveals patterns of resilience and vulnerability in discussions, such as those following disasters Randazzo et al. (2024). Temporal text networks Vega and Magnani (2018) and multimodal network models Bonifazi et al. (2023) further support analysis of communication across media and platforms.
@@
Multilayer approaches have also enriched opinion dynamics, where single-layer models often oversimplify how ideas form and spread. This approach provides a powerful framework for exploring how phenomena in one domain can influence behaviour in another. For example, a model using a communication layer for sharing opinions and an influence layer for observing actions, revealed a range of complex real-world outcomes Zino et al. (2020). These included rapid, widespread paradigm shifts, the establishment of unpopular norms where innovation adoption fails despite positive sentiment, and persistent community resistance to change.
@@
As discussed in the previous section, the interaction between different processes is a key area of study for multilayer network research. A recent work Amato et al. (2017) investigated a two-layered network, where agents held opinions on different issues in each layer. While each layer individually tended toward consensus, the interactions between layers could stabilize a mixed state where both opinions coexisted for a long period. This was most pronounced in sparse, positively correlated networks, which facilitated the formation of cross-layer opinion groups.
@@
The adaptive voter model Holme and Newman (2006) has been extended to multilayer networks to study how evolving connections affect collective opinion. While multiplexity (the fraction of common nodes across layers) can prevent complete fragmentation, it can also give rise to a new “shattered” fragmentation phase Diakonova et al. (2016). This phase is characterised by large consensus groups of opposite opinions coexisting with isolated nodes. A similar shattered fragmentation transition in the multiplex adaptive voter model with a notion of triadic closure has been used to describe the empirical distribution of cluster sizes in an online game Klimek et al. (2016b).
@@
Furthermore, a central question in this modelling paradigm is the extent to which such complex multilayer systems can be simplified, especially whether the dynamics on a multilayer network could be reduced to an equivalent single-layer representation. It turns out that conventional aggregation procedures fail to capture key non-linear effects, such as the prolonged lifetime of multiplex states at low interlayer connectivity Diakonova et al. (2016). This work suggests that there are fundamental, non-reducible differences between multilayer dynamics and their single-layer approximations, highlighting the necessity of explicitly modelling these intricate, interconnected structures to fully understand complex social phenomena.